if the lines joining the origin to the intersection points of the y=mx+2 and the curve x^2+y^2=1 are at right angle , then Plus 2 and the Cursed squared plus y squared

Grade:11

3 Answers

Piyush Kumar Behera
7 years ago
Make the question clear.I cant get what the question is trying to ask.The last part of 
the question is not  clear.
please use the latex to make it clear.
 
Prabhav Sharma
7 years ago
If the lines joining the origin to the intersecting points of the y=mx+2 and the curve x2+y2=1 are at right angle, then
(1) m2=1
(2) m2=3
(3) m2=7
(4)  2m2=1
now please give me the answer.cleared
 
Piyush Kumar Behera
7 years ago
There is a formula you can check it in any coordinate geometry book(such as s k goyal,etc)
which gives the joint equation of the pair of lines joining the origin and point of intersection of a curve and a straight line.
Joint equation of the lines joining the origin to the intersecting points of the y=mx+2 and the curve x2+y2=1
                                    x2+y2=((y-mx)/2)2
                                       =>       x2(4-m2)+2mxy+3y2=0
Since this lines are at right angles,
             => 4-m+3=0=>m2=7
so the ans is (3) m2=7
I hope you have understood,and if you haven’t then please read the above mentioned topic.Please upvote my solution to encourage me.